11. Use trig identities to solve the trig equation exactly for Osts 21. cos2t = -sin (t + /2)
Answer #1

criven that, cos (21) = -sin (t+1) > 2003?(H) - 1 = - cosct) [: COS(2+) = 2005%C4-17 - ano, sin (4+4)= cost -> 20052 (+) + C